\section*{References}

\begin{thebibliography}{99}
\bibitem{Vlec32}
	 J.\ H.\ Van Vleck,
	 The Theory of Electric and Magnetic Susceptibilities,
	 Oxford University Press, Oxford, 1932.
\bibitem{Coul56} 
	 C.\ A.\ Coulson and M.\ J.\ Stephen,
	 Proceedings of the Physical Society A 69  (1956) 777.
\bibitem{Turb87} 
	 A.\ V.\ Turbiner,
	 Yadernaya Fizika 46 (1987) 204 (in Russian).
\bibitem{Pote01} 
	 A.\ Y.\ Potekhin and A.\ V.\ Turbiner,  
	 Physical Review A 63 (2001) 065402. 
\bibitem{Szmy12}
   R. Szmytkowski, P. Stefa{\'n}ska,
   Physical Review A 85 (2012) 042502.
\bibitem{Szmy97}
	 R. Szmytkowski,
   Journal of Physics B 30 (1997) 825  
   [erratum: Journal of Physics B 30 (1997) 2747;
   addendum: arXiv:physics/9902050].
\bibitem{Szmy02b}
   R. Szmytkowski,
   Journal of Physics B 35 (2002) 1379. 
\bibitem{Szmy04}
   R. Szmytkowski, K. Mielewczyk,
   Journal of Physics B 37 (2004) 3961. 
\bibitem{Szmy02a}
   R. Szmytkowski,
   Physical Review A 65 (2002) 012503. 
\bibitem{Miel06}
   K. Mielewczyk, R. Szmytkowski,
   Physical Review A 73 (2006) 022511  
   [erratum:  Physical Review A 73 (2006) 039908].
\bibitem{Szmy11}
   R. Szmytkowski, P. Stefa{\'n}ska,
   e-print arXiv:1102.1811.
\bibitem{Stef12}
   P. Stefa{\'n}ska, R. Szmytkowski,
   International Journal of Quantum Chemistry 112 (2012) 1363. 
\bibitem{Szmy14}
   R. Szmytkowski and P. Stefa{\'n}ska, 
   Physical Review A 89 (2014) 012501.
\bibitem{Szmy16}
	 R. Szmytkowski and G. {\L}ukasik,
	 e-print arXiv:1602.00613.
\bibitem{Szmy16b}
	 R. Szmytkowski and G. {\L}ukasik,
	 Atomic Data and Nuclear Data Tables (2016), http://dx.doi.org/10.1016/j.adt.2016.02.002.
\bibitem{Stef16}
	 P. Stefa{\'n}ska, 
	 Physical Review A 93 (2016) 022504.
\bibitem{Stef15}
	 P. Stefa{\'n}ska, 
	 Physical Review A 92 (2015) 032504.
\bibitem{Stef16b}
	 P. Stefa{\'n}ska,
	 Atomic Data and Nuclear Data Tables 108 (2016) 193.
\bibitem{Mohr14}
	 P. J. Mohr, D. B. Newell, B. N. Taylor, 
	 CODATA Recommended Values of the Fundamental Physical Constants: 2014, arXiv:abs/1507.07956.


	

\end{thebibliography}

\documentclass[reviewcopy]{elsarticle}

\usepackage[reviewcopy]{adndt}
\usepackage{longtable}

\usepackage{amsmath}

\usepackage{amssymb}

\usepackage{lscape}

\biboptions{square,sort&compress}
\bibpunct[]{[}{]}{,}{n}{}{;}
\begin{document}

\begin{frontmatter}

\journal{Atomic Data and Nuclear Data Tables}


\title{Magnetic-dipole-to-electric-quadrupole cross-susceptibilities for relativistic hydrogenlike atoms in some low-lying discrete energy eigenstates}

  \author{Patrycja Stefa{\'n}ska\corref{cor1}}
  \ead{pstefanska@mif.pg.gda.pl}

  \cortext[cor1]{Corresponding author.}

  \address{Atomic Physics Division,
Department of Atomic, Molecular and Optical Physics, 
Faculty of Applied Physics and Mathematics, \\
Gda{\'n}sk University of Technology, 
Narutowicza 11/12, 80--233 Gda{\'n}sk, Poland}

\date{April 18, 2016} 


\begin{abstract}  
In this paper we present tabulated data for magnetic-dipole-to-electric-quadrupole cross-susceptibilities ($\chi_{\textrm{M}1 \to \textrm{E}2}$) for Dirac one-electron atoms with a pointlike, spinless and motionless nucleus of charge $Ze$. Numerical values of this susceptibility for the hydrogen atom ($Z=1$) and for hydrogenic ions with $2 \leqslant Z \leqslant 137$ are computed from the general analytical formula, recently derived by us [P. Stefa{\'n}ska, Phys. Rev. A 93 (2016) 022504], valid for an arbitrary discrete energy eigenstate.  In this work we provide 30 tables with the values of  $\chi_{\textrm{M}1 \to \textrm{E}2}$  for the ground state, and also for the first, the second and the third set of excited states (i.e.: 2s$_{1/2}$, 2p$_{1/2}$, 2p$_{3/2}$, 3s$_{1/2}$, 3p$_{1/2}$, 3p$_{3/2}$, 3d$_{3/2}$, 3d$_{5/2}$, 4s$_{1/2}$, 4p$_{1/2}$, 4p$_{3/2}$, 4d$_{3/2}$, 4d$_{5/2}$, 4f$_{5/2}$ and 4f$_{7/2}$) of the relativistic hydrogenlike atoms.  The value of the inverse of the fine-structure constant used in the calculations is $\alpha^{-1}=137.035 \: 999 \: 139$, and was taken from CODATA 2014.\\
\end{abstract}

\end{frontmatter}

\textbf{Keywords:} Hydrogenlike atom, Electromagnetic moments, Quadrupole moment, Magnetic field.

\begin{center}
\large{\textbf{Published as: At. Data Nucl. Data Tables 113 (2017) 316--349}
\\*[1ex]
\textbf{doi: 10.1016/j.adt.2016.04.004}\\*[5ex]}
\end{center}



\newpage

\tableofcontents
\listofDtables
\listofDfigures
\vskip5pc


\section{Introduction}

Theoretical investigation of atomic and molecular electromagnetic susceptibilities has been very commonly reported from early decades of the last century \cite{Vlec32}. Most often, numerical computation of such quantities for many-electron atoms and molecules, as well as analytical calculations for simple systems, can be compared with experimental data, and sometimes they appear to be very helpful in the design of other experiments or improving those already used. Thus, in this work we provide theoretical study of the magnetic-dipole-to-electric-quadrupole cross-susceptibility for hydrogenlike atoms with a pointlike, spinless and infinitely heavy nucleus of charge $Ze$. 

The phenomenon of inducing electric quadrupole moment $\mathcal{Q}$ in such systems by a perturbing uniform (dipole) magnetic field of a strength $B$ was pointed out already in the mid-1950s by Coulson and Stephen \cite{Coul56}. Analytical derivation of the expression for this moment was reported also later by Potekhin and Turbiner \cite{Turb87, Pote01}. However, in all these articles, calculations have been performed within the framework of the \emph{nonrelativistic} quantum mechanics. First fully \emph{relativistic} determination of that quantity has been carried out by us in 2012 \cite{Szmy12}. Using the perturbation theory combined with the technique based on the Sturmian expansion of the generalized Dirac--Coulomb Green function \cite{Szmy97}, we have successfully derived the analytical expression for $\mathcal{Q}^{(1)}(B)$ (the first-order correction to the electric quadrupole moment as a function of the field strength $B$) induced in the \emph{ground} state of Dirac one-electron atom (for other applications of this powerful technique to dipole electromagnetic properties of such system, see Refs. \cite{Szmy02b,Szmy04,Szmy02a,Miel06,Szmy11,Stef12,Szmy14}).

Very recently, Szmytkowski and {\L}ukasik considered analytically \cite{Szmy16} and numerically \cite{Szmy16b} some families of \emph{multipole} electric and magnetic susceptibilities for the ground state of relativistic hydrogenlike atoms. In particular, they pointed out that $2^{\lambda}$-pole electric moments induced in the system by static $2^{L}$-pole magnetic field are closely related to $2^{L}$-pole magnetic moments induced in the atom by static $2^{\lambda}$-pole electric field. To be precise, they indicated that the values of the magnetic-to-electric ($\chi_{\textrm{M}L \to \textrm{E}\lambda}$) and electric-to-magnetic ($\alpha_{\textrm{E}\lambda \to \textrm{M}L}$) cross-susceptibilities are \emph{equal} for ions with the charge numbers from the whole range $1 \leqslant Z \leqslant 137$. This suggests that theoretical knowledge of such kind of quantities -- in particular, the $\chi_{\textrm{M}1 \to \textrm{E}2}$ susceptibility -- may be doubly useful for experimentalists: those who would like to measure one of these atomic parameters but are unable to do that, may try to conduct another experiment which allows to determine the second of them.

Some time ago, we have discussed the aforementioned physical quantity in terms of another generalization. In \cite{Stef16} we derived  analytically the expression for the magnetic-dipole-to-electric-quadrupole cross-susceptibility ($\chi_{\textrm{M}1 \to \textrm{E}2}$) for the atom being in an \emph{arbitrary} discrete energy eigenstate, characterized by the set of quantum numbers $\{n_{r}, \kappa, \mu \}$, in which $n_{r}$ denotes the radial quantum number, the Dirac quantum number $\kappa$ is an integer different form zero, whereas $\mu=-|\kappa|+\frac{1}{2}, -|\kappa|+\frac{3}{2}, \ldots, |\kappa|-\frac{1}{2}$ is the magnetic quantum number. The calculations for such general case were based on the Sturmian expansion of the generalized Dirac--Coulomb Green function \cite{Szmy97} (for other applications of this method to another magnetic dipole properties of excited atomic states, see Refs. \cite{Stef15,Stef16b}). In view of the definition below:
\begin{equation}
\mathcal{Q}^{(1)}=\textrm{sgn}(\mu)  (4\pi \epsilon_0) c \: \chi_{\textrm{M}1 \to \textrm{E}2} B,
\label{eq:1}
\end{equation}
where $c$ is the speed of light, while $\mathcal{Q}^{(1)} \equiv \mathcal{Q}_{20}^{(1)}$ is the only one of the spherical components $\mathcal{Q}_{2M}^{(1)}$ ($M=0,\pm1,\pm2$) of the quadrupole electric moment tensor which can be induced in the atom by the perturbing magnetic field $\boldsymbol{B}=B\boldsymbol{n}_z$ (for details, see Ref.\ \cite{Stef16}),  the formula for the atomic parameter under study can be written as follows:
\begin{eqnarray}
\chi_{\textrm{M}1 \to \textrm{E}2}(n_{r},\kappa,\mu)
&=&
\frac{\alpha a_0^4}{Z^4}  
\frac{|\mu|}{64(4\kappa^2-1)^2} 
\Bigg\{
\Theta_{n_{r}\kappa \mu}^{(\texttt{I})}
+\sum_{\substack{\kappa'=-\infty \\ (\kappa' \neq 0)}}^{\infty} 
\frac{\eta_{\kappa \mu}^{(+)}\delta_{\kappa',-\kappa+1}+\eta_{\kappa \mu}^{(-)}\delta_{\kappa',-\kappa-1}}{N_{n_{r}\kappa}+\kappa'} 
\nonumber \\
&& \times
\Bigg[
\Theta_{n_{r}\kappa}^{(\texttt{II})}
+\frac{3n_{r}!N_{n_{r}\kappa}^3\Gamma(n_{r}+2\gamma_{\kappa}+1)}{(N_{n_{r}\kappa}-\kappa)(\gamma_{\kappa'}-\gamma_{\kappa}-n_{r}+1)\Gamma(2\gamma_{\kappa'}+1)} 
\sum_{k=0}^{n_{r}}\sum_{p=0}^{n_{r}} \mathcal{X}(k) \mathcal{Y}(p) 
\nonumber \\
&& \quad \times 
 {}_3F_2 \left(\begin{array}{c} 
\gamma_{\kappa'}-\gamma_{\kappa}-k-2,\:
\gamma_{\kappa'}-\gamma_{\kappa}-p-1,\: 
\gamma_{\kappa'}-\gamma_{\kappa}-n_{r}+1 \\
\gamma_{\kappa'}-\gamma_{\kappa}-n_{r}+2,\:
2\gamma_{\kappa'}+1
\end{array}
;1 
\right)
\Bigg]
\Bigg\}.
\label{eq:2}
\end{eqnarray}
Here $\alpha$ is the Sommerfeld's fine structure constant, $a_0$ denotes, as usually, the Bohr radius, $\Gamma(\zeta)$ is the Euler's gamma function and ${}_3F_2$ is the generalized hypergeometric function, while
\begin{equation}
\eta_{\kappa \mu}^{(\pm)}=\frac{2\kappa \pm 1}{2\kappa \mp 3}\left[(2\kappa \mp 1)^2-4\mu^2\right],
\label{eq:3}
\end{equation} 
\begin{eqnarray}
\Theta_{n_{r}\kappa \mu}^{(\texttt{I})}=32 \left(\alpha Z\right)^2\kappa^2(4\kappa^2-12\mu^2-1) 
\left\{(5n_{r}^2+10n_{r} \gamma_{\kappa}+4\gamma_{\kappa}^2+1)N_{n_{r}\kappa}^2-\kappa(n_{r}+\gamma_{\kappa})\left[2\kappa(n_{r}+\gamma_{\kappa})+3N_{n_{r} \kappa}\right] \right.\},
\label{eq:4}
\end{eqnarray}
\begin{eqnarray}
\Theta_{n_{r}\kappa}^{(\texttt{II})}=6(N_{n_{r}\kappa}-\kappa)N_{n_{r}\kappa}^2 \left\{
4\kappa (n_{r}+\gamma_{\kappa})^2\left[7(n_{r}+\gamma_{\kappa})^2-3\gamma_{\kappa}^2+5\right]+N_{n_{r}\kappa}\left[63(n_{r}+\gamma_{\kappa})^4+70(n_{r}+\gamma_{\kappa})^3\right. \right.
\nonumber \\
\left. \left. -21(2\gamma_{\kappa}^2-5)(n_{r}+\gamma_{\kappa})^2-10(3\gamma_{\kappa}^2-5)(n_{r}+\gamma_{\kappa})+3(\gamma_{\kappa}^2-1)(\gamma_{\kappa}^2-4) \right] 
\right\},
\label{eq:5}
\end{eqnarray}
\begin{equation}
\mathcal{X}(k)=\frac{(-)^{k}}{k!(n_{r}-k)!} \frac{\Gamma(\gamma_{\kappa}+\gamma_{\kappa'}+k+3)}{\Gamma(k+2\gamma_{\kappa}+1)}\left[C_{k}^{(1)}(N_{n_{r}\kappa}+\kappa')-C_{k}^{(2)}(n_{r}-k-3) \right],
\label{eq:6}
\end{equation}
\begin{equation}
\mathcal{Y}(p)=\frac{(-)^{p}}{p!(n_{r}-p)!} \frac{\Gamma(\gamma_{\kappa}+\gamma_{\kappa'}+p+2)}{\Gamma(p+2\gamma_{\kappa}+1)}\left[(n_{r}-p)(\kappa+\kappa')+2(N_{n_{r}\kappa}-\kappa) \right],
\label{eq:7}
\end{equation}
with
\begin{equation}
C_{k}^{(1)}=(n_{r}-k)+\frac{n_{r}+\gamma_{\kappa}}{N_{n_{r}\kappa}}\left(\kappa-N_{n_{r}\kappa}\right), \qquad  \qquad C_{k}^{(2)}=\frac{n_{r}+\gamma_{\kappa}}{N_{n_{r}\kappa}}(n_{r}-k)+\left(\kappa-N_{n_{r}\kappa}\right),
\label{eq:8}
\end{equation}
\begin{equation}
N_{n_{r}\kappa}=\sqrt{n_{r}^2+2n_{r}\gamma_{\kappa}+\kappa^2}
\label{eq:9}
\end{equation}
and
\begin{equation}
\gamma_{\kappa}=\sqrt{\kappa^2-(\alpha Z)^2}
\label{eq:10}.
\end{equation}
In Ref.\ \cite{Stef16} we have shown that the expression from Eq.\ (\ref{eq:2}) remains valid for an \emph{arbitrary} discrete energy eigenstate of the atom. Nevertheless, we realize that its quite complexity form might be perceived by someone as inconvenient for calculations. Such our thoughts, as well as the lack of any tabulated data in the aforementioned article, were the main reasons for which we decided to perform numerical calculations of that quantity. In the next Section we will discuss briefly our results.

\section{Discussion of results}

The computational results obtained with the aid of the formula in Eq.\ (\ref{eq:2}) are presented here in the form of 30 tables comprising the values of the $\chi_{\textrm{M}1 \to \textrm{E}2}$  for the relativistic hydrogenlike atoms (with the nuclear charge number from the range $1 \leqslant Z \leqslant 137$) being in the ground state 1s$_{1/2}$, or in any state belonging to the first three sets of excited states, i.e.: 2s$_{1/2}$, 2p$_{1/2}$, 2p$_{3/2}$, 3s$_{1/2}$, 3p$_{1/2}$, 3p$_{3/2}$, 3d$_{3/2}$, 3d$_{5/2}$, 4s$_{1/2}$, 4p$_{1/2}$, 4p$_{3/2}$, 4d$_{3/2}$, 4d$_{5/2}$, 4f$_{5/2}$ and 4f$_{7/2}$, having regarded all possible values of the magnetic quantum number $\mu$. 

All tabulated data were obtained with the use of the current value $\alpha^{-1}=137.035 \: 999 \: 139$ of the inverse of the fine-structure constant recommended by the Committee on Data for Science and Technology (CODATA 2014) \cite{Mohr14}.

Values for the susceptibility in question for the atomic ground state (Table \ref{tab:1s_1-2}) fully coincide with those obtained by Szmytkowski and {\L}ukasik \cite{Szmy16b}. Analysis of all numbers from Tables \ref{tab:1s_1-2}--\ref{tab:4f_7-2_mu_7-2} allows one to conclude that the sign of $\chi_{\textrm{M}1 \to \textrm{E}2}$ is related to the sign of the angular-plus-parity symmetry quantum number $\kappa$, i.e.: generally, if $\kappa>0$ then $\chi_{\textrm{M}1 \to \textrm{E}2}$ takes the positive values, and if $\kappa<0$ then  $\chi_{\textrm{M}1 \to \textrm{E}2}$ is negative. Exceptions to this rule are states 1s$_{1/2}$, 2s$_{1/2}$ and those of the type $n$p$_{3/2}$, with $|\mu|=1/2$. For these latter there is the critical value of nuclear charge number $Z$, $Z_c$, such that for $Z<Z_c$ the $\chi_{\textrm{M}1 \to \textrm{E}2}$ susceptibility is negative, and for $Z \geqslant Z_c$ it takes the positive values. For the state 2p$_{3/2}$ it equals  $Z_c=123$, and increases for the higher excited states, i.e. for the 3p$_{3/2}$ state there is $Z_c=127$, and for the state 4p$_{3/2}$ we have $Z_c=129$. Furthermore, for states with orbital angular momentum $l=2$ (the d-type), the values of $\chi_{\textrm{M}1 \to \textrm{E}2}$ for every ions with $1 \leqslant Z \leqslant 137$, calculated  with $|\mu|=3/2$, are approximately two times greater than appropriate numbers obtained with the use of $|\mu|=1/2$. Moreover, the susceptibility under study for d-type states with total angular momentum $j=3/2$ takes nearly the same values as those with $j=5/2$, but with the opposite sign (this is caused by opposite directed spins). In turn, in determination of $\chi_{\textrm{M}1 \to \textrm{E}2}$ for the f-type states, there is no matter, which magnetic quantum number is taken into account: differences between values obtained with $|\mu|=3/2$ and those with $|\mu|=5/2$ are of the order of $10^{-8}$ in the units of $a_0^{4}$.



\ack
I would like to thank Professor A.\ Rutkowski for his strong suggestion to publish present results. Useful discussions with Professor R.\ Szmytkowski are also acknowledged.



\clearpage

\newpage

\TableExplanation

In all the tables we have used the following notation:

\bigskip
\renewcommand{\arraystretch}{1.0}



\end{document}